A military aircraft flying at an altitude of 1,000 m has launched a glide missile. If the missile landed 100m from the launch point, how fast was the aircraft flying?

1
Expert's answer
2022-05-02T13:41:29-0400

The horizontal range

"R=v_0\sqrt{2H/g}"

Hence, the speed of the aircraft

"v_0=R/\sqrt{2H/g}\\
=100/\sqrt{2*1000/9.8}=7\:\rm m/s"